hi i'm having an extreme amount of trouble setting up word problems for linear equations , especially this one .

Your teacher is giving you a test worth 100 points containing 40 questions. There are two-point and four-point questions on the test. How many of each type of question are on the test?

Found 2 solutions by stanbon, fgwjr2003:
Answer by stanbon(75887)

Your teacher is giving you a test worth 100 points containing 40 questions. There are two-point and four-point questions on the test. How many of each type of question are on the test?
--------------------------------------
Equations:
t + f = 40
2t+4f = 100
-----------------
Modify:
t + f = 40
t + 2f = 50
-------------------
Subtract 1st from 2nd to solve for "f":
f = 10 (# of 4 point questions)
---
Substitute to solve for "t"
t + 10 = 40
t = 30 (# of 2 point questions)
===================================
Cheers,
Stan H.

Answer by fgwjr2003(2)

number of questions___+___value of question___=______total value of question
________x_________+__________2____________=_________2x_______________
______40-x________x__________4____________=_______4(40-x)_____________
_____________________________total # of questions_=_____100______________
2x + 4(40-x)=100
2x + 160 - 4x = 100
2x-4x
-2x + 160 = 100
____-100_-100
-2x + 60 = 0
____-60_-60
-2x/-2 = -60/-2
__x____=___30
therfore 30-2 pt questions
40-30 = 10-4pt questions